E5.6 (LO 1), AP PCB Corporation manufactures a single product. Monthly production costs incurred in process are shown below for the production of 3,000 units. The utilities and mainte- manu nance costs are mixed costs. The fixed portions of these costs are $300 and $200, respectively Production in Units Production Costs Direct materials Direct labor Utilities Property taxes Indirect labor Supervisory salaries Maintenance 7,500 18,000 2.100 1,000 1,900 1,100 2,400 Instructions a. Identify the above costs as variable, fixed, or mixed. b. Calculate the expected costs when production is 5,000 units.
Answer #1

a.

Production Cost

Direct material

Variable cost

Direct labor

Variable cost

Utilities

Mixed cost

Property taxes

Fixed cost

Indirect labor

Variable cost

Supervisory salaries

Fixed cost

Maintenance

Mixed cost

Depreciation

Fixed cost

b.

Fixed cost = Property taxes + Supervisory salaries + Depreciation + Fixed utilities + Fixed maintenance

                = $ 1,000 + 1,900 + $ 2,400 + $ 300 + $ 200 = $ 5,800

Variable cost for 3,000 units = Direct material + Direct labor + Indirect labor

                                         = $ 7,500 + $ 18,000 + $ 4,500 = $ 30,000

Variable cost per unit = $ 30,000/3,000 = $ 10

Variable cost portion of mixed cost = Total cost – fixed portion

Variable utilities cost for 3,000 units = $ 2,100 - $ 300 = $ 1,800

Variable utilities cost per unit = $ 1,800/3,000 = $ 0.60

Variable maintenance cost for 3,000 units = $ 1,100 - $ 200 = $ 900

Variable maintenance cost per unit = $ 900/3,000 = $ 0.30

Total variable cost per unit = $ 10 + $ 0.60 + $ 0.30 = $ 10.90

Expected cost for 5,000 units = Monthly fixed cost + Variable cost per unit x 5,000

                                          = $ 5,800 + $ 10.90 x 5,000

                                            = $ 5,800 + $ 54,500 = $ 60,300

Expected cost is $ 60,300 when production is 5,000 units